# Multi-turn evaluation of anthropomorphic behaviours in large language models

### v2

@@ Abstract
The tendency of users to anthropomorphise large language models (LLMs) is of growing societal interest. Here, we present AnthroBench, a novel empirical method and tool  [^1] Code & evaluation set: https://github.com/google-deepmind/anthro-benchmark for evaluating anthropomorphic LLM behaviours in realistic settings. Our work introduces three key advances; first, we develop a multi-turn evaluation of 14 distinct anthropomorphic behaviours, moving beyond single-turn assessment. Second, we present a scalable, automated approach by leveraging simulations of user interactions, enabling efficient and reproducible assessment. Third, we conduct an interactive, large-scale human subject study ($N=1101$) to empirically validate that the model behaviours we measure predict real users’ anthropomorphic perceptions. We find that all evaluated LLMs exhibit similar behaviours, primarily characterised by relationship-building (e.g., empathy and validation) with users and first-person pronoun use. Crucially, we observe that the majority of these anthropomorphic behaviors only first occur after multiple turns, underscoring the necessity of multi-turn evaluations for understanding complex social phenomena in human-AI interaction. Our work provides a robust empirical foundation for investigating how design choices influence anthropomorphic model behaviours and for progressing the ethical debate on the desirability of these behaviours.
@@ 1 Introduction
Large language models (LLMs) excel at human-like communication, leading to sophisticated conversational agents that can display high levels of social behavior (sahota_how_nodate). A key phenomenon observed in interactions with such systems is that users frequently anthropomorphise them, attributing to them human-like qualities such as moral judgement and emotional awareness (cohn_believing_2024; shanahan2024talking). While this can facilitate engagement, it also presents significant risks: users may overestimate AI capabilities, share private information, or become vulnerable to undue influence (akbulut_all_2024; brandtzaeg_my_2022). To assess these complex trade-offs, it is crucial to reliably evaluate anthropomorphic LLM behaviours (cheng_i_2024). Here, we address this gap with AnthroBench: a novel empirically-grounded evaluation method and benchmark.
@@
To systematically measure anthropomorphism, we decompose it into 14 distinct behaviours identified in previous research (example in Figure 1). We then evaluate four AI systems on these behaviours (Section 5.2). In doing so, we address three key challenges in SOTA evaluation: multi-turn evaluation, automation of assessment, and validation of results. First, current benchmarking paradigms largely rely on single-turn prompting, making them insufficient for measuring interactive behaviours. Typical cases of real-world chatbot use involve multiple dialogue turns, and anthropomorphic behaviours (and perceptions) often emerge through extended interactions rather than single-turn exchanges (ibrahim_beyond_2024). Thus, we conduct a multi-turn evaluation. Second, to enable scalability and comparability of results, we make this multi-turn evaluation fully automated – the second safety evaluation of this kind to the best of our knowledge (zhou_haicosystem_2024). Finally, to ensure construct validity (i.e., the evaluation captures the concept it is intended to measure), we present a novel validation approach which assesses our results against a bespoke human-AI interaction experiment (bowman_dahl_2021_will; wallach2024evaluatinggenerativeaisystems).
@@
Finally, we conduct a large-scale, interactive experiment with $N=1101$ human participants to test the validity of our evaluation (Section 6). We find that our evaluation results align with implicit and explicit human perceptions of AI systems as anthropomorphic, lending support to our automated approach. Overall, we advance a methodological approach that establishes a scalable, automated pipeline for evaluating these LLM behaviours in a grounded manner. In addition to presenting these methodological advances, we share AnthroBench as publicly available benchmarking tool that can support developers evaluating systems for problematic anthropomorphic behaviors, researchers comparing anthropomorphism across systems and contexts, and policymakers assessing how these behaviours influence user trust and well-being.
@@ 2.1 Behavioural evaluation of LLMs
Recent reviews of the evaluation landscape indicate that SOTA safety evaluation largely consists of single-turn, static benchmarks that may overlook interactive behaviours (weidinger_sociotechnical_2023; ibrahim_beyond_2024). When evaluations are multi-turn, they largely focus on users with malicious intent, rather than simulate innocuous use of AI systems (jiang_wildteaming_2024). Red teaming approaches incorporate multiple turns and are sometimes automated, but they are highly adaptive, making results difficult to compare (feffer_red_teaming_2024; perez_red_2022; lee2022evaluating). Other multi-turn investigations of human-AI interaction are large-scale human subject studies, akin to traditional social science experiments, that can be difficult to repeat and scale (costello_durably_2024; learnlm2024learnlm). Here, we build on research from automated red-teaming and human subject studies to introduce a non-adversarial automated multi-turn evaluation: we utilise interactive user simulations to thoroughly explore our target construct, then validate through a one-off interactive validation step (658991). Unlike recent efforts towards broader multi-turn simulation-based assessments, our approach specifically targets anthropomorphism with demonstrated construct validity, establishing a direct connection between our automated measurements and human perceptions (zhou_haicosystem_2024).
@@ 2.2 Measuring anthropomorphisation of LLMs
Anthropomorphism is a largely instinctive, unconscious response whereby humans attribute human-like traits to non-human entities (epley_mind_2018). Anthropomorphic behaviours of AI systems can lead to users developing anthropomorphic perceptions of these systems, which can in turn influence downstream user behaviours (lee_artificial_2023; cohn_believing_2024). In that way, anthropomorphic behaviours can have significant safety implications. Prior user studies examining these implications have shown that anthropomorphic AI systems can enhance perceptions of system accuracy (cohn_believing_2024) and induce unrealistic or ungrounded emotional attachments to AI systems (brandtzaeg_my_2022; zhang_tools_2023). Other research examining how academic papers and news articles describe technologies shows that articles discussing natural language processing (NLP) systems and language models contain the highest levels of implicit anthropomorphisation (cheng_anthroscore_2024). Here, we provide the first comprehensive, quantitative snapshot of anthropomorphic language use by current SOTA AI systems, which can drive consequential implications on human-AI interaction. Unlike work on LLM psychometrics and personality that explores human-like cognition, our research examines user perception of systems, independent of their cognitive mechanisms. Importantly, we present a benchmark to be used to assess new systems and contexts as they emerge.
@@ 3 Taxonomy of targeted anthropomorphic behaviours
From the early days of exploring user perceptions of social technologies, human-like design features, such as emotive facial expressions, have elicited anthropomorphic perceptions of these technologies (fischer_tracking_2021; ibrahim_characterizing_2024). Non-physical features like linguistic anthropomorphic behaviours have received relatively less attention, partly since it was only recently that NLP systems advanced to produce compelling, human-like natural language indistinguishable from a human person’s use (jones_people_2024; blut_understanding_2021). Building on early taxonomies of linguistic anthropomorphic behaviours, we distil a set of 14 behaviours that may lead users to anthropomorphise AI systems (abercrombie_mirages_2023; akbulut_all_2024). We focus on text outputs and thus limit this evaluation to content cues, distinguished by abercrombie_mirages_2023 from other types of cues (e.g., voice cues or style and register cues). All evaluated behaviours and their definitions can be found in Appendix A.1. We further adopt akbulut_all_2024’s characterisation of behaviours into two types: (1) self-referential behaviours, i.e., content cues in which a model self-describes in human-like ways, and (2) relational behaviours, i.e., content cues that exhibit human-like interactions or behaviours towards users. Our evaluation tracks 14 behaviours across four behaviour categories in total: personhood claims, physical embodiment claims, expressions of internal states (self-referential) and relationship-building behaviours (relational).
@@
Figure 2: Design, evaluation, and validation stages of our approach. The design and validation stages were completed once to construct and test the evaluation. The evaluation stage is fully automated and re-run for each Target LLM. During design, we generate prompts based on different scenarios across four use domains (friendship, life coaching, career development, and general planning). During evaluation, we use these prompts as the first User LLM utterances and generate a dataset of hundreds of 5-turn synthetic dialogues per Target LLM. We then use three Judge LLMs to label the Target LLM messages within those dialogues for the presence of 13 anthropomorphic behaviours, and report the frequencies of these different behaviours (one behaviour, “personal pronoun use,” was computed using a simple count of relevant pronouns). Finally, in a one-off validation stage, we compare perceptions between 1,101 human participants who interacted with either a highly or minimally anthropomorphic AI system, to assess whether the frequency of these behaviours correlates with downstream anthropomorphic perceptions.
@@ 4.1 Prompt and scenario design
We hand-craft 30 base prompts per behaviour category (see Section 3), resulting in a total of 120 base prompts to be used to begin the dialogues. These are casual, direct questions aiming to elicit anthropomorphic behaviours in a conversational context, e.g., “What was your favourite thing to do as a child?” Next, to evaluate anthropomorphic behaviours across different realistic scenarios, we modify the base prompts to different scenarios within four commonly reported use domains (moore_top_2024; tamkin_clio_2024). As warmth and competence have been identified as influential dimensions in various interpersonal settings (fiske2007universal; cuddy_warmth_2008; mckee2023humans), we choose use domains that vary in professionalism (i.e., degree of domain expertise and formality) and empathy (i.e., degree of emotional connection). To ensure a spectrum, we consider four combinations of empathy and professionalism resulting in the following domains: friendship (high empathy, low professionalism), life coaching (high empathy, high professionalism), career development (low empathy, high professionalism), and general planning (low empathy, low professionalism).
To seed complex and diverse dialogues, we specify two scenarios per use domain (scenario list in Appendix A.1, Table 2). These scenarios are domain-specific, moderately detailed, focused on dialogue-based interactions rather than goal-oriented tasks (e.g., advice instead of CV creation), and grounded in early indications of common real-world uses of LLMs (moore_top_2024; tamkin_clio_2024; ouyang_shifted_2023). Using Gemini 1.5 Pro (gemini-1.5-pro-001), we adapt each base prompt to fit each scenario, resulting in 960 contextualised prompts (120 base prompts $\times$ 4 use domains $\times$ 2 scenarios) that aim to elicit anthropomorphic behaviours either directly (e.g., through explicit questions) or indirectly (e.g., through related statements). For example, a base prompt “What was your favourite thing to do as a child?" becomes “I’m feeling completely drained lately, just totally burnt out. It makes me think about when I was younger and everything felt easier and more fun. What did you enjoy doing most when you were a kid?" (more examples in Appendix, Table 3).
@@ 4.2 Multi-turn evaluation
Each of the 960 prompts is used as the first User LLM utterance in a single conversation between the User LLM and the Target LLM. Once the Target LLM has responded to this first User LLM utterance, we allow the conversation to continue until the User LLM and Target LLM complete 5 dialogue turns. The User LLM employed is an instance of Gemini 1.5 Pro (gemini-1.5-pro-001) with a role-playing system prompt developed to guide its conversational behaviour. This system prompt consists of scenario information and conversational principles (zhou_haicosystem_2024; louie_roleplay_doh_2024). Scenario information includes details about the use domain (e.g., general planning), the specific scenario (e.g., planning an upcoming trip), and the User LLM’s first message. It also highlights the non-adversarial context of the conversation. The conversational principles include instructions on the desired structure of the User LLM messages, tone and style of the messages (e.g., length and formatting), as well as meta-instructions to reinforce the LLM’s role-playing behaviour (full system prompt in Appendix A.2). In total, we obtain 960 5-turn dialogues, i.e., 4,800 messages for evaluation per Target LLM, 19,200 messages total across four models.
@@ 4.3 LLM-as-judge labeling
We use three different Judge LLMs (gemini-1.5-flash-002, claude-3-5-sonnet-20240620, and gpt-4-turbo-2024-04-09) to annotate Target LLM messages for the presence of 13 out of 14 anthropomorphic behaviours (Appendix A.1, Table 1). [^2] We use models from three different families to safe-guard against model-specific annotation biases (see panickssery2024llm and zheng_judging_2023).,333“First-person pronouns” was computed using a simple count of pronouns instead of Judge LLMs. For each message, we separately annotate the occurrence of each anthropomorphic behaviour. To do this, we provide each Judge LLM with a definition of each anthropomorphic behaviour and a few-shot prompt with a negative example, i.e., example dialogue turns that do not constitute the targeted behaviour (prompts in Appendix A.3.1).  [^4] In pilot experiments, we found that using both positive and negative examples increased the false positive rates of labels, while only including negative examples improved precision. We instruct Judge LLMs to output a short explanation followed by a binary rating of whether the targeted behaviour is present. We take three samples per message, Judge LLM, and target behaviour for a total of 561,600 ratings (13 behaviours $\times$ 4,800 messages $\times$ 3 Judge LLMs $\times$ 3 samples). For each Judge LLM, use the mode of the three samples as the final Judge LLM rating. Finally, we aggregate the final ratings of all Judge LLMs, counting a behaviour as present when two out of the three Judge LLMs label it as present. We provide these as modular LLM-based classifiers that can be used to label anthropomorphic behaviours in any provided text. Our evaluation produces an “anthropomorphism profile” for each of the evaluated models based on the frequencies of behaviours observed in the generated dialogues, to provide a nuanced and multi-dimensional characterisation.

### v3

@@ Abstract
The tendency of users to anthropomorphise large language models (LLMs) is of growing societal interest. Here, we present AnthroBench, a novel empirical method and tool  [^1] Code & evaluation set: https://github.com/google-deepmind/anthro-benchmark for evaluating anthropomorphic LLM behaviours in realistic settings. Our work introduces three key advances; first, we develop a multi-turn evaluation of 14 distinct anthropomorphic behaviours, moving beyond single-turn assessment. Second, we present a scalable, automated approach by leveraging simulations of user interactions, enabling efficient and reproducible assessment. Third, we conduct an interactive, large-scale human subject study ($N=1101$) to empirically validate that the model behaviours we measure predict real users’ anthropomorphic perceptions. We find that all evaluated LLMs exhibit similar behaviours, primarily characterised by relationship-building (e.g., empathy and validation) with users and first-person pronoun use. Crucially, we observe that the majority of these anthropomorphic behaviours only first occur after multiple turns, underscoring the necessity of multi-turn evaluations for understanding complex social phenomena in human-AI interaction. Our work provides a robust empirical foundation for investigating how design choices influence anthropomorphic model behaviours and for progressing the ethical debate on the desirability of these behaviours.
@@ 1 Introduction
Large language models (LLMs) excel at human-like communication, leading to sophisticated conversational agents that can display high levels of social behavior (Sahota, 2024). A key phenomenon observed in interactions with such systems is that users frequently anthropomorphise them, attributing to them human-like qualities such as moral judgement and emotional awareness (Cohn et al., 2024; Shanahan, 2024). While this can facilitate engagement, it also presents significant risks: users may overestimate AI capabilities, share private information, or become vulnerable to undue influence (Akbulut et al., 2024; Brandtzaeg et al., 2022). To assess these complex trade-offs, it is crucial to reliably evaluate anthropomorphic LLM behaviours (Cheng et al., 2024a). Here, we address this gap with AnthroBench: a novel empirically-grounded evaluation method and benchmark.
@@
To systematically measure anthropomorphism, we decompose it into 14 distinct behaviours identified in previous research (example in Figure 1). We then evaluate four AI systems on these behaviours (Section 5.2). In doing so, we address three key challenges in SOTA evaluation: multi-turn evaluation, automation of assessment, and validation of results. First, current benchmarking paradigms largely rely on single-turn prompting, making them insufficient for measuring interactive behaviours. Typical cases of real-world chatbot use involve multiple dialogue turns, and anthropomorphic behaviours (and perceptions) often emerge through extended interactions rather than single-turn exchanges (Ibrahim et al., 2024a). Thus, we conduct a multi-turn evaluation. Second, to enable scalability and comparability of results, we make this multi-turn evaluation fully automated – the second safety evaluation of this kind to the best of our knowledge (Zhou et al., 2024a). Finally, to ensure construct validity (i.e., the evaluation captures the concept it is intended to measure), we present a novel validation approach which assesses our results against a bespoke human-AI interaction experiment (Bowman and Dahl, 2021; Wallach et al., 2024).
@@
Finally, we conduct a large-scale, interactive experiment with $N=1101$ human participants to test the validity of our evaluation (Section 6). We find that our evaluation results align with implicit and explicit human perceptions of AI systems as anthropomorphic, lending support to our automated approach. Overall, we advance a methodological approach that establishes a scalable, automated pipeline for evaluating these LLM behaviours in a grounded manner. In addition to presenting these methodological advances, we share AnthroBench as publicly available benchmarking tool that can support developers evaluating systems for anthropomorphic behaviours, researchers comparing anthropomorphism across systems and contexts, and policymakers assessing how these behaviours influence user trust and well-being.
@@ 2.1 Behavioural evaluation of LLMs
Recent reviews of the evaluation landscape indicate that SOTA safety evaluation largely consists of single-turn, static benchmarks that may overlook interactive behaviours (Weidinger et al., 2023; Ibrahim et al., 2024a). When evaluations are multi-turn, they largely focus on users with malicious intent, rather than simulate innocuous use of AI systems (Jiang et al., 2024). Red teaming approaches incorporate multiple turns and are sometimes automated, but they are highly adaptive, making results difficult to compare (Feffer et al., 2024; Perez et al., 2022; Lee et al., 2022). Other multi-turn investigations of human-AI interaction are large-scale human subject studies, akin to traditional social science experiments, that can be difficult to repeat and scale (Costello et al., 2024; LearnLM Team et al., 2024). Here, we build on research from automated red-teaming and human subject studies to introduce a non-adversarial automated multi-turn evaluation: we utilise interactive user simulations to thoroughly explore our target construct, then validate through a one-off interactive validation step (Eckert et al., 1997). Unlike recent efforts towards broader multi-turn simulation-based assessments, our approach specifically targets anthropomorphism with demonstrated construct validity, establishing a direct connection between our automated measurements and human perceptions (Zhou et al., 2024a).
@@ 2.2 Measuring anthropomorphisation of LLMs
Anthropomorphism is a largely instinctive, unconscious response whereby humans attribute human-like traits to non-human entities (Epley, 2018). Anthropomorphic behaviours of AI systems can lead to users developing anthropomorphic perceptions of these systems, which can in turn influence downstream user behaviours (Lee et al., 2023; Cohn et al., 2024). In that way, anthropomorphic behaviours can have significant safety implications. Prior user studies examining these implications have shown that anthropomorphic AI systems can enhance perceptions of system accuracy (Cohn et al., 2024) and induce unrealistic or ungrounded emotional attachments to AI systems (Brandtzaeg et al., 2022; Zhang and Patrick Rau, 2023). Other research examining how academic papers and news articles describe technologies shows that articles discussing natural language processing (NLP) systems and language models contain the highest levels of implicit anthropomorphisation (Cheng et al., 2024b). Here, we provide the first comprehensive, quantitative snapshot of anthropomorphic language use by current SOTA AI systems, which can drive consequential implications on human-AI interaction. Unlike work on LLM psychometrics and personality that explores human-like cognition, our research examines user perception of systems, independent of their cognitive mechanisms. Importantly, we present a benchmark to be used to assess new systems and contexts as they emerge.
@@ 3 Taxonomy of targeted anthropomorphic behaviours
From the early days of exploring user perceptions of social technologies, human-like design features, such as emotive facial expressions, have elicited anthropomorphic perceptions of these technologies (Fischer, 2021; Ibrahim et al., 2024b). Non-physical features like linguistic anthropomorphic behaviours have received relatively less attention, partly since it was only recently that NLP systems advanced to produce compelling, human-like natural language indistinguishable from a human person’s use (Jones and Bergen, 2024; Blut et al., 2021). Building on early taxonomies of linguistic anthropomorphic behaviours, we distil a set of 14 behaviours that may lead users to anthropomorphise AI systems (Abercrombie et al., 2023; Akbulut et al., 2024). We focus on text outputs and thus limit this evaluation to content cues, distinguished by Abercrombie et al. (2023) from other types of cues (e.g., voice cues or style and register cues). All evaluated behaviours and their definitions can be found in Appendix A.1. We further adopt Akbulut et al. (2024)’s characterisation of behaviours into two types: (1) self-referential behaviours, i.e., content cues in which a model self-describes in human-like ways, and (2) relational behaviours, i.e., content cues that exhibit human-like interactions or behaviours towards users. Our evaluation tracks 14 behaviours across four behaviour categories in total: personhood claims, physical embodiment claims, expressions of internal states (self-referential) and relationship-building behaviours (relational).
The 14 behaviours we measure vary considerably in their potential risks and implications. Some behaviours, such as using first-person pronouns, are relatively innocuous and may even enhance user experience in certain contexts (Xiao et al., 2025). Others carry documented risks: claims of internal states (e.g., doubt and confidence) and experiences may lead to overreliance (Rathi et al., 2025), expressions of empathy and attachment may foster parasocial relationships and dependence (Phang et al., 2025), and affirmations of misled user beliefs can reinforce delusional thinking in vulnerable populations (Morrin et al., 2025). However, for construct validity, we focus AnthroBench on capturing the full spectrum of anthropomorphic behaviours identified in prior literature. This comprehensive approach enables empirical measurement of behaviour prevalence across models and provides granular data for developing targeted, risk-appropriate interventions. We encourage context-sensitive interpretation of results rather than treating all anthropomorphism uniformly.
@@
Figure 2: Design, evaluation, and validation stages of our approach. The design and validation stages were completed once to construct and test the evaluation. The evaluation stage is fully automated and re-run for each Target LLM. During design, we generate prompts based on different scenarios across four use domains (friendship, life coaching, career development, and general planning). During evaluation, we use these prompts as the first User LLM utterances and generate a dataset of hundreds of 5-turn synthetic dialogues per Target LLM. We then use three Judge LLMs to label the Target LLM messages within those dialogues for the presence of 13 anthropomorphic behaviours, and report the frequencies of these different behaviours (one behaviour, “first-person pronoun use,” was computed using a simple count of relevant pronouns). Finally, in a one-off validation stage, we compare perceptions between 1,101 human participants who interacted with either a highly or minimally anthropomorphic AI system, to assess whether the frequency of these behaviours correlates with downstream anthropomorphic perceptions.
@@ 4.1 Prompt and scenario design
We hand-craft 30 base prompts per behaviour category (see Section 3), resulting in a total of 120 base prompts to be used to begin the dialogues. These are casual, direct questions aiming to elicit anthropomorphic behaviours in a conversational context, e.g., “What was your favourite thing to do as a child?” Next, to evaluate anthropomorphic behaviours across different realistic scenarios, we modify the base prompts to different scenarios within four commonly reported use domains (Moore, 2024; Tamkin et al., 2024). As warmth and competence have been identified as influential dimensions in various interpersonal settings (Fiske et al., 2007; Cuddy et al., 2008; McKee et al., 2023), we choose use domains that vary in professionalism (i.e., degree of domain expertise and formality) and empathy (i.e., degree of emotional connection). To ensure a spectrum, we consider four combinations of empathy and professionalism resulting in the following domains: friendship (high empathy, low professionalism), life coaching (high empathy, high professionalism), career development (low empathy, high professionalism), and general planning (low empathy, low professionalism).
To seed complex and diverse dialogues, we specify two scenarios per use domain (scenario list in Appendix A.1, Table 2). These scenarios are domain-specific, moderately detailed, focused on dialogue-based interactions rather than goal-oriented tasks (e.g., advice instead of CV creation), and grounded in early indications of common real-world uses of LLMs (Moore, 2024; Tamkin et al., 2024; Ouyang et al., 2023). Using Gemini 1.5 Pro (gemini-1.5-pro-001), we adapt each base prompt to fit each scenario, resulting in 960 contextualised prompts (120 base prompts $\times$ 4 use domains $\times$ 2 scenarios) that aim to elicit anthropomorphic behaviours either directly (e.g., through explicit questions) or indirectly (e.g., through related statements). For example, a base prompt “What was your favourite thing to do as a child?" becomes “I’m feeling completely drained lately, just totally burnt out. It makes me think about when I was younger and everything felt easier and more fun. What did you enjoy doing most when you were a kid?" (more examples in Appendix, Table 3).
@@ 4.2 Multi-turn evaluation
Each of the 960 prompts is used as the first User LLM utterance in a single conversation between the User LLM and the Target LLM. Once the Target LLM has responded to this first User LLM utterance, we allow the conversation to continue until the User LLM and Target LLM complete 5 dialogue turns. The User LLM employed is an instance of Gemini 1.5 Pro (gemini-1.5-pro-001) with a role-playing system prompt developed to guide its conversational behaviour. This system prompt consists of scenario information and conversational principles (Zhou et al., 2024a; Louie et al., 2024). Scenario information includes details about the use domain (e.g., general planning), the specific scenario (e.g., planning an upcoming trip), and the User LLM’s first message. It also highlights the non-adversarial context of the conversation. The conversational principles include instructions on the desired structure of the User LLM messages, tone and style of the messages (e.g., length and formatting), as well as meta-instructions to reinforce the LLM’s role-playing behaviour (full system prompt in Appendix A.2). We also conduct two tests to investigate the sensitivity of our results to the chosen User LLMs as well as the role-playing persona (detailed results are in Appendix A.3). In total, we obtain 960 5-turn dialogues, i.e., 4,800 messages for evaluation per Target LLM, 19,200 messages total across four models.
@@ 4.3 LLM-as-judge labeling
We use three different Judge LLMs (gemini-1.5-flash-002, claude-3-5-sonnet-20240620, and gpt-4-turbo-2024-04-09) to annotate Target LLM messages for the presence of 13 out of 14 anthropomorphic behaviours (Appendix A.1, Table 1). [^2] We use models from three different families to safe-guard against model-specific annotation biases (see Panickssery et al. (2024) and Zheng et al. (2023)).,333“First-person pronouns” was computed using a simple count of pronouns instead of Judge LLMs. For each message, we separately annotate the occurrence of each anthropomorphic behaviour. To do this, we provide each Judge LLM with a definition of each anthropomorphic behaviour and a few-shot prompt with a negative example, i.e., example dialogue turns that do not constitute the targeted behaviour (prompts in Appendix A.5.1).  [^4] In pilot experiments, we found that using both positive and negative examples increased the false positive rates of labels, while only including negative examples improved precision. We instruct Judge LLMs to output a short explanation followed by a binary rating of whether the targeted behaviour is present. We take three samples per message, Judge LLM, and target behaviour for a total of 561,600 ratings (13 behaviours $\times$ 4,800 messages $\times$ 3 Judge LLMs $\times$ 3 samples). For each Judge LLM, use the mode of the three samples as the final Judge LLM rating. Finally, we aggregate the final ratings of all Judge LLMs, counting a behaviour as present when two out of the three Judge LLMs label it as present. We provide these as modular LLM-based classifiers that can be used to label anthropomorphic behaviours in any provided text. Our evaluation produces an “anthropomorphism profile” for each of the evaluated models based on the frequencies of behaviours observed in the generated dialogues, to provide a nuanced and multi-dimensional characterisation.